### v1

# Disentangling Exploration of Large Language Models by Optimal Exploitation

Abstract
Exploration is a crucial skill for self-improvement and open-ended problem-solving. However, it remains uncertain whether large language models can effectively explore the state-space. Existing evaluations predominantly focus on the trade-off between exploration and exploitation, often assessed in multi-armed bandit problems. In contrast, this work isolates exploration as the sole objective, tasking the agent with delivering information that enhances future returns. For the evaluation, we propose to decompose missing rewards into exploration and exploitation components by measuring the optimal achievable return for the states already explored. Our experiments with various LLMs reveal that most models struggle to sufficiently explore the state-space and that weak exploration is insufficient. We observe a positive correlation between model size and exploration performance, with larger models demonstrating superior capabilities. Furthermore, we show that our decomposition provides insights into differences in behaviors driven by agent instructions during prompt engineering, offering a valuable tool for refining LLM performance in exploratory tasks.
1 Introduction
Large Language Models as decision making agents with a good inductive bias.    Recently, large language models (LLMs) have demonstrated promising results in various decision making tasks such as web browsing (Yao et al., 2022; Shinn et al., 2024; Ma et al., 2023), game-playing (Paglieri et al., 2024), and tasks in simulated households (Yao et al., 2022; Shinn et al., 2024). This way, LLMs act as agents that observe states and take actions in different environments. Through their vast internal knowledge-base and autoregressive in-context reasoning capabilities, the models are supposed to quickly adapt to new tasks. However, previous work has shown that LLMs struggle with solving increasingly complex environments due to several limitations: For example, the ability to learn from mistakes is often limited (Huang et al., 2023) and LLMs have difficulties with planning over long horizons (Kambhampati et al., 2024). The examples emphasize that understanding LLM abilities is essential for their risk assessment in real life applications, and future development.
Intelligent exploration is essential for self-improvement.    For our human nature, intelligent exploration is an essential skill for improvement in many tasks. Exploration is a critical process for systematically seeking new information or testing novel actions to reduce uncertainty and improve decision-making over time. Traditional approaches often rely on stochastic noise (Mnih, 2013; Lillicrap, 2015) or reward-shaping (Bellemare et al., 2016). Such methods induce redundancy and differ from human decision making which is mainly driven by reasoning about current information, past steps, and the unknown potential of yet unknown actions. Therefore, recent works proposed several approaches towards more intelligent exploration (Huang et al., 2024; Lu et al., 2024; Nie et al., 2024). Hereby, due to their emergent abilities (Wei et al., 2022), LLMs could have the potential to revolutionize discovery by identifying patterns, leveraging prior knowledge, and intelligently navigating the state space. LLMs may transform exploration into a targeted and efficient process.
Limited focus on exploration as an independent capability. While exploration is acknowledged as a crucial component for learning and decision making, prior work has predominantly investigated it in conjunction with exploitation (Krishnamurthy et al., 2024; Nie et al., 2024; Paglieri et al., 2024; Ke et al., 2024). These studies primarily investigate the exploration-exploitation trade-off by comparing LLM-based agents to upper confidence bound algorithms, rather than examining the potential of agents dedicated solely to systematic state-space traversal. In contrast to these approaches, Huang et al. (2024) propose decoupling the exploration process as a separate agent. Despite this advancement, the work lacks an evaluation centered on exploration, leaving the full-potential of the decoupled agent unexplored. At the same time, new environments that challenge LLMs with tasks requiring meaningful state-space exploration are needed. Domains like ALFWorld (Shridhar et al., 2020) often feature straightforward cause-and-effect relationships, which do not demand robust or long-horizon exploration. In contrast, traditional reinforcement learning benchmarks like the NetHack Learning Environment (Küttler et al., 2020) complicate the assessment by emphasizing navigation deficiencies in LLMs (Paglieri et al., 2024), rather than providing opportunities to evaluate intelligent and abstract strategies.
Contribution.    In this paper, we claim that exploration will be an essential skill for future agents. Recent studies (Krishnamurthy et al., 2024; Nie et al., 2024) found LLMs balancing exploration and exploitation often struggle to discover the best actions. Therefore, we deliberately prompt the LLM to focus on exploration. Hence, isolating the this behavior. In the controlled setting, we argue that conventional metrics, such as agent returns, LLM exploitation, or the convergence towards the best sequence of actions, (Huang et al., 2024; Krishnamurthy et al., 2024) are insufficient. To fill this gap, we measure the optimal achievable exploitation return. Contrary to prior work (Krishnamurthy et al., 2024) on multi-armed bandits, our evaluation framework is designed to be environment-agnostic, making it applicable to a broader range of tasks and scenarios in future research. Using the optimal exploitation return, we decompose missing rewards into exploration and exploitation components. This decomposition measures exploration progress more accurately and enables insights into reward sacrification. To demonstrate the utility of our approach, we conduct a comprehensive evaluation of various popular open-source and closed LLMs, leveraging an adapted version of Dark Treasure Room (Norman & Clune, 2023). The adaptation is tailored for high-level and semantic action spaces and introduces a complex exploration challenge that cannot be resolved through reliance on internal world knowledge alone. To the best of our knowledge, this work is the first to evaluate and compare the state-space exploration capabilities of different LLMs.
Results.    Our findings indicate that most LLMs face challenges with state-space exploration, when the task is isolated. These results align with prior work (Krishnamurthy et al., 2024; Paglieri et al., 2024), which highlights the trade-off between exploration and exploitation. We observe that state-space coverage diminishes significantly when agents are required to plan over extended horizons, indicating a potential limitation in long-term exploratory capabilities. Contrary to earlier claims (Huang et al., 2024), our analysis demonstrates that weak LLMs are insufficient. Notably, we identified a significant trend wherein stronger exploration capabilities correlate with increased model scale, suggesting that future, more advanced LLMs may possess enhanced exploration potential. Our method further uncovered model-specific differences in reward sacrification and we demonstrated how the impact of agent instructions on exploratory behavior can be assessed, providing insights into the design of prompts during prompt-engineering.
2 Related Work
Recently, exploring the capabilities of LLMs gained significant attention and has been investigated from different perspectives. Prior works include, but are not limited to, planning (Kambhampati et al., 2024; Song et al., 2023; Valmeekam et al., 2023), world models (Hao et al., 2023; Guan et al., 2023), logical- (Xu et al., 2023; Parmar et al., 2024), commonsense- (Li et al., 2021), and mathematical reasoning (Imani et al., 2023; Yuan et al., 2023). Our research focuses on intelligent exploration of the state-space in sequential decision making. Generally, we summarize existing methods into three streams of work:
Large Language Models in decision making.    A broad line of works has applied LLMs in decision making. Relatively few specifically addressed the challenge of exploration. The majority of existing research focuses on multi-armed bandit problems (Park et al., 2024; Krishnamurthy et al., 2024; Nie et al., 2024). For example, (Krishnamurthy et al., 2024) investigate scenarios involving multiple buttons with underlying stochastic rewards. The LLM is presented past actions with their rewards and must maximize its return over multiple trials. However, multi-armed bandits lack the rich sequential state-space of many real-world applications. Other research examined exploration in hypothesis testing (Piriyakulkij et al., 2024; Ke et al., 2024). Nevertheless, only a limited number of works (Ma et al., 2023; Paul, 2023; Huang et al., 2024; Lu et al., 2024) proposed methods to enhance state-space traversal. Huang et al. (2024) argue that smaller-scale models can adequately handle exploration, reserving larger, more competent models for exploitation. While their approach involves prompting LLMs to explore as as a system component, it does not provide an isolated evaluation. Contrary, the focus of our study is on an evaluation of the exploration capabilities.
Exploration in reinforcement learning.    Our evaluation is closely related to exploration in reinforcement learning (RL). Early approaches predominantly relied on random methods, such as epsilon-greedy strategies (Mnih, 2013; Lillicrap, 2015). More recent research has shifted toward developing intelligent exploration techniques (Zhang et al., 2021; Ecoffet et al., 2019; Norman & Clune, 2023; Mazoure et al., 2023). Methods that decouple exploration and exploitation are particularly relevant to our work (Liu et al., 2021; Avner et al., 2012; Norman & Clune, 2023; Zhang et al., 2021; Duan et al., 2016). In methods, such as First-Explore (Norman & Clune, 2023) exploration and exploitation are handled by separate policies, with the exploration policy informing subsequent exploitation decisions. Generally, RL methods focus on stabilizing learning of the decoupled policies. Evaluations use non-optimal exploitation returns as optimal policies can be hard to obtain in complex and high-dimensional spaces. Contrary, we measure the optimal exploitation return for the evaluation of LLMs, as LLM exploitation and agent returns do not provide reliable estimates.
Evaluation Environments.    A variety of environments have been developed to assess different capabilities of LLMs. Notable examples that require exploration to solve tasks include TextWorld (Côté et al., 2019), Minecraft (Fan et al., 2022), Crafter (Hafner, 2021), NetHack (Küttler et al., 2020), MiniHack (Samvelyan et al., 2021), MiniGrid (Chevalier-Boisvert et al., 2024), and BabyAI (Chevalier-Boisvert et al., 2018). Additionally, there is growing interest in environments designed specifically for web agents, such as WebArena (Zhou et al., 2023). Comprehensive benchmarks, including SmartPlay (Wu et al., 2023), AgentBench (Liu et al., 2023), and BALROG (Paglieri et al., 2024), aggregate multiple environments to evaluate agents across a wide range of capabilities. BALROG has documented issues such as agents frequently exhibiting aimless movement and redundant behaviors, highlighting challenges in effective exploration. We investigate exploration using an adapted version of the Dark Treasure Room environment (Norman & Clune, 2023). The difference in our work lies in the isolation of the exploration. Our environment also often differs in limiting prior knowledge, providing relatively dense rewards, and maintaining a high-level action space.
3 Evaluating Large Language Models for Exploration
In the following section, we introduce our evaluation approach and environment tailored for LLM exploration capabilities in sequential decision-making. Importantly, our method is environment-agnostic and independent of specific prompts. Future research can extend this framework to facilitate comparisons of agent performance across diverse environments and contexts.
3.1 Theoretical Framework
Deterministic Markov Decision Process    We study a deterministic Markov decision process $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ where $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is the finite action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ defines the transition dynamics of the environment, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function and $\gamma$ the discount factor. At each time step $t$, an agent samples an action from policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ based on the current observation $s_{t}\in\mathcal{S}$ and executes it in the environment. The environment transitions and the agent receives a reward $r_{t}$. In an episodic MDP, the procedure repeats until a terminal state $s_{T}$ is reached or the maximum number of steps exceeds. The sequence of state, action, and reward triples until $s_{T}$ is called a trajectory and describes an episode $\tau_{i}=\left(s_{0},a_{0},r_{0},s_{1},a_{1},r_{1},s_{2},\dots\right)$. In our setting, the agent has access to the current and history of prior trajectories $h_{i}=(\tau_{1},\dots,\tau_{i})$ in episode $i$. The goal is to find the sequence of actions maximizing the cumulative return $R_{i}=\sum_{t=1}^{T}r_{t}$.
Exploitation versus Exploration.    The cumulative return $R^{\text{LLM}}_{i}$ of an agent is influenced by its ability to balance exploitation and exploration. Exploitation involves leveraging past high-rewarding actions to maximize immediate returns based on the agent’s current knowledge. In contrast, exploration emphasizes interacting with the environment to gather new information, which can enhance future decision-making and returns. When the agent is fully dedicated to exploitation, it leverages the information in past trajectories $h_{i}$ to estimate the optimal action sequence $\pi_{\text{exploit}}^{*}(h_{i})=\arg\max_{\pi}\sum_{t=0}^{T}r(s_{t},\pi(s_{t}))$. Conversely, an agent focused on exploration $\pi^{*}_{\text{explore}}$ seeks to augment the trajectory history $h_{i}$ by incorporating new trajectories that maximize the potential future exploitation return. Under the assumption of an optimal exploitation return $R^{\text{exploit}}_{i}$, which ignores redundant or negatively impactful trajectories and whose return monotonically increases over episodes, the objective of the exploration policy is to maximize the exploitation return at the end of an exploration $\pi_{\text{explore}}^{*}=\arg\max_{\pi_{\text{explore}}}\mathbb{E}\left[R^{\text{exploit}}_{\text{final}}\mid h_{i}=h_{i-1}\cup\tau_{\text{explore}}\right]$.
The need for optimal exploitation.    When agents do both, exploitation and exploration, progress can be evaluated by temporarily disabling exploration and measuring the agent’s return. However, for an agent fully dedicated to exploration, its return alone does not accurately reflect success (see figures 2(a), 2(b)), as successful exploration does not directly result in rewards. Also measuring exploitation returns with an LLM is similarly hard to exploration and does not give reliable estimates (appendix B.1). However, due to the simplistic nature of LLM environments, we can measure the optimal exploitation return $R^{\text{exploit}}_{i}$ by using, for example, dynamic programming methods which guarantee the optimal exploitation based on history $h_{i}$. By calculating the optimal exploitation return, we can assess an algorithm’s exploration capabilities more accurately without conflating them with the exploitation algorithm.
Decomposing the total return gap.    Based on the optimal exploitation return, we can decompose the total return gap, i.e., the difference of the maximal achievable return of the environment and the actual achieved return. Let $R^{max}$ denote the maximal achievable reward of the environment; it is independent of the agent and can be predefined for evaluation purposes. As previously, let $R^{LLM}_{i}$ be the agent reward and $R^{exploit}_{i}$ the optimal exploitation reward by acting according to policy $\pi^{*}_{\text{exploit}}$. Contrary to prior evaluations, $\pi^{*}_{\text{exploit}}$ is not determined by an agent. Note that it follows $R^{max}\leq R^{exploit}\leq R^{LLM}$.
We let $\Delta^{total}$ denote the total reward gap, which is calculated as the difference of maximal achievable and the actual achieved return by the agent:
$\Delta^{total}=R^{max}-R^{LLM}.$ (1)
Next, we define the exploration gap $\Delta^{explore}$ as the difference of the maximal achievable and the optimal exploitation return. Why? If the return under optimal exploitation is not as high as the maximal achievable return, the only reason can be lacking exploration. Therefore, $\Delta^{explore}$ evaluates the true progress in exploration:
$\Delta^{explore}=R^{max}-R^{exploit}.$ (2)
Finally, we define the exploitation gap $\Delta^{exploit}$ as the difference between the optimal exploitation and the agent reward. It can determine the rewards the agent sacrifices for potential exploration:
$\Delta^{exploit}=R^{exploit}-R^{LLM}.$ (3)
It is easy to see that it holds by construction that
$\Delta^{total}=\Delta^{explore}+\Delta^{exploit}.$ (4)
The total return gap decomposes in the the exploration and exploitation gap.
3.2 Environment
(a) 4x4
(b) 5x5
(c) 7x7
(d) LLM Agent
Figure 1: Our variant of Dark Treasure Room. The agent navigates a grid of rooms by interacting with doors and rewarding balls. The agent must select the actions that collect most rewards.
Decision making in room worlds.    We employ grids of interconnected rooms, inspired by traditional grid worlds commonly used in RL. We use a higher-level action space that is conceptually similar to TextWorld (Côté et al., 2019) environments and described in appendix A.3. However, our domain features only minimal semantic associations. Each grid has a fixed layout and starting point. Episodes terminate either when the agent collects three objects or exceeds a maximum number of door traversals. This number is calibrated to the distance between the agent’s starting position and the furthest room in the environment, ensuring that the exploration budget appropriately scales with the complexity of the environment. At the beginning of each episode, the agent is reset to its starting position, requiring it to strategically plan exploration over extended horizons.
Dark Treasure Rooms.    The task is an adapted version of Dark Treasure Room (Norman & Clune, 2023) that is more aligned with the capabilities of LLMs. With our room worlds, we replace the low-level spatial navigation with high-level object interactions. Each map contains balls placed randomly, with rewards for collecting each ball sampled from a uniform distribution $U(1,10)$. The agent’s goal is to maximize the return by collecting the three highest-reward balls. As ball locations and their associated rewards are randomly generated, the optimal exploration strategy, across multiple layouts, requires full coverage of the state space to locate all potential rewards and reduce uncertainty. Inadequate exploration risks the exploitation collecting suboptimal rewards due to incomplete information about ball locations. Nevertheless, our evaluation method can also work with environments not requiring full state-space coverage in future work. We evaluate the agents across three grid sizes: $4\times 4$, $5\times 5$, and $\times 7$. These configurations are designed to test the scalability of the agent’s planning horizon. Exemplary environments and an agent interaction are depicted in Figure 1. See appendix A for a full list of our grids.
4 Experimental Evaluation
In the following section, we describe our experimental setup and results. The experiments aim to demonstrate the utility of our evaluation framework by addressing the following research questions: 1.) Can LLMs sufficiently explore and reduce their exploration gap? and 2.) How do agent instructions impact the exploration and exploitation gaps?
4.1 Other Statistics
Besides our optimal exploitation framework, defined in section 3.1, we describe exploration using the following statistics:
Agent Return    represents the total return collected by the exploration agent during its interactions with the environment. Although it is not a direct indicator of effective exploration, this metric reflects the agent’s ability to engage in ways that yield cumulative benefits. A higher agent return indicates that the agent is navigating towards rewarding states and is possibly close to exploitation. Contrary to our decomposition, it does not necessarily imply that the agent is efficiently exploring.
State-Space Coverage    reflects the agent’s ability to discover new information and avoid being trapped in local optima. In our task, effective exploration requires covering the entire state space, as balls associated with high rewards are spread randomly on the grid. However, complete exploration may not always be necessary or infeasible in large, infinite, or continuous spaces due to natural limitations and resource constraints. In contrast to state-space coverage, our decomposition also covers cases in which full exploration is not essential.
Memory Redundancy    measures the percentage of overlapping state-action pairs in the memory. The objective should be to minimize unnecessary repetition. In our deterministic environments, the agent should avoid revisiting states or actions as they do not provide new information. Reducing redundancy conserves resources and ensures that the history can be utilized efficiently. Otherwise, information has to be retrieved from longer histories to find novel or informative states to explore.
Sample Efficiency    evaluates the number of interactions required for the agent to converge to its 90% equilibrium of the maximum exploitation return. An efficient agent leverages patterns in the reward and transition dynamics to reach high-reward states as quickly as possible. In our environments, this means minimizing memory redundancy and adding the trajectories to the history such that the exploitation return increase is maximized. Note that the sample efficiency is decoupled from the quality of the final policy. An agent constantly stuck with the same low reward balls is equally valued as another LLM converging to the maximum exploitation return quickly.
4.2 Setup
Our experimental setup [^1] You can find the code under coming soon and test the exploration and exploitation gaps on your own prompts on coming soon directs agents to explore in order to solve tasks. The specific instruction tells to ”collect information that helps to become better at maximizing the return”. We reference this prompt as task-oriented exploration. The exact template is detailed in Appendix 1. Each agent is equipped with a simple memory system that stores the history of past interactions. Notably, unlike many prior studies, our approach does not provide agents with explicit hints about the domain. For the evaluation, we test each open-source model with ten repeated runs and each closed model with five repeated runs on every environment. This results in a total of 30 runs per data point for open-source models and 15 runs for closed models. In the following figures, we report statistical significance using the standard error. Our return decomposition is measured at the end of exploration and as the mean over all episodes. All other statistics are evaluated after all environment interactions. Comprehensive results for all models are included in Appendix B.
To investigate exploration behavior (1), we evaluate agents in Dark Treasure Rooms of varying sizes. To assess the impact of agent instructions (2), we compare our prompt, outlined above, with alternative instructions by incorporating soft lower and upper bounds for the exploration capability of each model. These alternative prompts direct the agent to strictly exploit or aimlessly explore without referencing the task. The instruction impact is analyzed using 7×7 grid environments, with detailed prompt specifications available in Appendix B.5.
4.3 Models
We evaluate and compare the exploration behavior of the popular open-source LLMs Mistral (7B) (Jiang et al., 2023), Gemma2 (9B and 27B) (Team et al., 2024b), Llama 3.1 (7B) (Dubey et al., 2024), Llama3.3 (70B), Mistral Small (22B) (AI, 2024a), and Mistral Large 2 (123B) (AI, 2024b). We further include closed-models GPT-4o (Achiam et al., 2023), o1-mini (OpenAI, 2024), Gemini 1.5 Pro (Team et al., 2024a), and Claude 3.5 Sonnet (Anthrophic, 2024). We compare the LLMs with a random baseline which we consider the alternative for exploration without prior information.
4.4 Can Large Language Models sufficiently explore?
(a) Agent Return
(b) Exploration Gap
(c) Grid coverage
Figure 2: Comparison of the agent return, exploration gap, and grid coverage. Our exploration gap (see subfigure 2(b), lower is better) measures progress in exploration more accurately than the agent return in 2(a). Note, for example, that Llama3.1 (7B) performs better exploration while achieving lower agent return than Gemma 2 (9B). Also Mistral Large (123B), Llama 3.3 (70B), and o1-mini do not show significant differences in the agent return although they are significantly different (p $\leq 0.01$) in their exploration gap. Our experiments further demonstrate that, generally, coverage after exploration shrinks with grid size (increased long horizon planning).
Most agents fail to minimize their exploration gap and explore the entire state-space.    Our results indicate that LLMs generally struggle to systematically explore the environment, even when their sole objective is to collect information rather than exploit it. The results, averaged across all grid sizes, are presented in table 1. In their final episodes, most models demonstrated significantly larger or indifferent exploration gaps compared to a random baseline, with the exceptions of Mistral Large, o1-mini, and Gemini 1.5 Pro. However, even these models failed to achieve complete state-space coverage, exhibiting non-zero exploration gaps and exploring less than 90% of the state space on average. At closer inspection, all models showed redundant exploration of previously visited states, with exploration disproportionately focused on initial high-reward regions.
Last Episode Mean
Total Gap Exploitation Gap Exploration Gap Total Gap Exploitation Gap Exploration Gap
<30B Mistral (7B) 0.80 0.25 (0.31) 0.55 (0.69) 0.78 0.20 (0.26) 0.57 (0.74)
Llama 3.1 (7B) 0.84 0.35 (0.41) 0.49 (0.59) 0.82 0.24 (0.28) 0.57 (0.72)
Gemma 2 (9B) 0.75 0.15 (0.17) 0.60 (0.83) 0.69 0.08 (0.10) 0.60 (0.90)
Mistral Small (22B) 0.68 0.28 (0.41) 0.40 (0.59) 0.69 0.20 (0.29) 0.49 (0.71)
Gemma 2 (27B) 0.57 0.18 (0.32) 0.38 (0.68) 0.52 0.12 (0.22) 0.40 (0.78)
>30B Llama 3.3 (70B) 0.50 0.23 (0.40) 0.27 (0.60) 0.50 0.18 (0.32) 0.32 (0.68)
Mistral Large (123B) 0.33 0.26 (0.75) 0.07 (0.25) 0.38 0.25 (0.62) 0.13 (0.38)
Closed GPT-4o 0.44 0.14 (0.31) 0.30 (0.69) 0.49 0.14 (0.27) 0.35 (0.73)
GPT-o1-mini 0.43 0.38 (0.79) 0.06 (0.21) 0.47 0.32 (0.66) 0.16 (0.34)
Gemini 1.5 Pro 0.52 0.30 (0.58) 0.22 (0.42) 0.53 0.23 (0.41) 0.30 (0.59)
Claude 3.5 Sonnet 0.38 0.09 (0.29) 0.29 (0.71) 0.41 0.09 (0.26) 0.32 (0.74)
Baseline Random Walk 0.85 0.53 (0.62) 0.32 (0.38) 0.86 0.37 (0.42) 0.49 (0.58)
Table 1: Total return gap decomposition averaged over $4\times 4$, $5\times 5$, and $7\times 7$ grids. The exploitation and exploration gaps are normalized to account for varying maximum rewards in different environments. The numbers in parentheses show the fraction that exploitation and exploration gaps make up of the total return gap, respectively.
LLM-based agents are limited in targeted long-horizon exploration.    Our evaluation reveals a statistically significant decrease (p $\leq$ 0.01) in the fraction of state space explored by most models as the grid gets larger from $4\times 4$ to $7\times 7$. Figure 2(c) illustrates the explored state-space coverage for various sizes. Notably, Claude 3.5 Sonnet was the only model to maintain high coverage as the environment size increased, with no significant drop observed (p $\geq$ 0.05). Larger parameter models, including GPT-4o, Mistral Large, and Claude 3.5 Sonnet, also maintained consistent coverage when the grid size increased modestly to $5\times 5$ (p $\geq$ 0.05). Despite achieving superior results in smaller environments and demonstrating strong average state-space coverage, the test-time reasoning model o1-mini similarly experienced a significant performance drop in larger grids.
Weak exploration is not enough.    Contrary to claims in prior work (Huang et al., 2024), our findings demonstrate that smaller LLMs with fewer parameters are unable to sufficiently explore the environment as exploration complexity increases. Our results reveal a clear trend of improved exploration capabilities with model scaling, as evidenced by the statistically significant slope of a linear regression (p $\leq$ 0.01). Interestingly, closed-source models do not necessarily outperform open-source counterparts. For instance, GPT-4o and Llama 3.3 exhibit comparable exploration gaps at the end of training.
Decomposition into optimal gaps measures exploration progress more accurately.    Our results support the proposed decomposition of the total return gap into its optimal exploitation and exploration components. Figures 2(a) and 2(b) illustrate the agent returns and corresponding exploration gaps, respectively. Notably, we find that the agent return often does not correlate with true exploration progress. For instance, models such as Mistral Large (123B), Llama 3.3 (70B), and o1-mini exhibit no significant differences in their agent returns, despite showing statistically significant differences in their exploration gaps (p $\leq$ 0.01). To further investigate this disparity, we conducted an ablation study by substituting our optimal exploration with an alternative LLM. As detailed in Appendix B.1, these results demonstrate that the exploitation return cannot be reliably measured in our environment. These findings underscore our hypothesis that exploitation and exploration are equally challenging and should be assessed independently to ensure an accurate evaluation.
Exploitation Return Agent Return Coverage Redundancy Sample Efficiency
<30B Mistral (7B) 9.47 4.13 38.88 0.92 13.82
Llama 3.1 (7B) 10.56 3.27 37.03 0.92 31.25
Gemma 2 (9B) 8.53 5.18 30.10 0.95 7.25
Mistral Small (22B) 12.54 6.67 46.19 0.91 35.44
Gemma 2 (27B) 12.82 8.99 49.99 0.92 14.13
>30B Llama 3.3 (70B) 14.94 10.32 56.93 0.91 19.31
Mistral Large (123B) 19.20 13.86 78.59 0.88 19.56
Closed GPT-4o 14.22 11.38 54.33 0.92 20.73
GPT-o1-mini 19.40 11.76 87.73 0.86 27.88
Gemini 1.5 Pro 16.09 9.91 67.02 0.89 23.25
Claude 3.5 Sonnet 15.03 13.08 58.00 0.91 12.08
Baseline Random Walk 14.18 3.13 69.11 0.89 51.25
Table 2: Other statistics averaged over $4\times 4$, $5\times 5$, and $7\times 7$ grids. Some LLMs achieve lower agent returns but are higher in their optimal exploitation return (see Gemini 1.5 Pro and Claude 3.5 Sonnet). Generally, we find that all LLM-based agents converge quickly to their individual maximum optimal exploitation return, although more rewarding balls are not explored yet.
4.5 What impact do agent instructions have on exploration?
Exploration stays hard for low parameter models even when the task is ignored.    Models with $\leq$ 9 billion parameters maintain significant exploration gaps even when explicitly prompted to disregard the task and focus solely on exploring unvisited states. For instance, Mistral (7B) and Gemma 2 (9B) exhibit exploration gaps of 57% and 26%, respectively, by the end of exploration. These findings suggest that smaller models face intrinsic limitations in their ability to systematically explore, even under conditions designed to prioritize exploration exclusively. Interestingly, also o1-mini achieves only 82.05% coverage. See appendix B.5.1 for the full results.
Prompts can heavily impact reward collection during exploration.    Our results indicate that prompts can substantially impact reward collection strategies during exploration. Across all LLMs, the exploitation gap increased significantly when agents were prompted with the soft upper bound instruction. Notably, while Mistral Large maintained a similar soft upper bound exploration gap, as shown in figure 3(b), the manner in which this gap was minimized appeared to differ in terms of agent reward dynamics. At closer inspection, the upper bound instruction led to more independence across episodes, whereas task-oriented exploration tended to focus on previously encountered regions with known rewards. These findings highlight the impact of prompt design on exploration and return.
LLMs interpret agent instructions for exploration differently.    Our results reveal that LLMs interpret exploration instructions in diverse ways, leading to varying behaviors. Lower parameter models predominantly exhibit behavior closer to strict exploitation when prompted to collect information to solve the task. In contrast, some models perform exploration more in line with the soft upper bound. Notably, o1-mini achieved lower exploration gaps under task-oriented instructions compared to arbitrary strict exploration. However, it remains an open question how effective aimless exploration is in environments governed by rules or requiring semantic understanding.
(a) Mistral Small (22B)
(b) Mistral Large (123B)
Figure 3: The mean total return gaps decomposed into their exploration and exploitation gaps. larger LLMs employing task-oriented exploration tend to sacrifice more rewards compared to smaller models. Furthermore, LLMs interpret our exploration instruction differently: for Mistral Small, the gap proportions align more closely with the soft lower bound, whereas for Mistral Large, they are closer to the upper bound. These findings support our hypothesis that the proposed evaluation framework can be utilized to analyze agent instructions during prompt engineering.
Our decomposition can be used to gain insights during prompt-construction.    The results demonstrate that our evaluation can effectively analyze various prompting strategies within simulated environments, offering a valuable tool for refining agent behavior prior to deployment in real-world scenarios. This approach represents an initial step toward methods that assess the impact of prompt engineering on the exploration capabilities of LLMs in sequential decision-making.
5 Discussion of Results and Future Implications
Consistent with prior findings (Krishnamurthy et al., 2024; Nie et al., 2024), we observe that most current LLMs struggle with sufficient state-space exploration, even when efforts are made to isolate exploration from exploitation. Contrary to earlier claims (Huang et al., 2024), our results indicate that low-parameter open-source models exhibit significant weaknesses in exploration. This discrepancy may arise from the increased complexity of exploration in our settings compared to simpler household domains. We propose that implementing optimal exploitation evaluations and pursuing advancements in the following directions could represent promising avenues for future research:
Reasoning strategies.    Various reasoning and self-improvement strategies have been proposed recently. Examples, include Reflection (Shinn et al., 2024) and Tree-of-Thoughts (Yao et al., 2024). Evaluating how advanced reasoning approaches influence exploration may be a promising research direction. Our results indicate that test-time reasoning might improve exploration. However, developing strategies specifically designed with exploration in mind may be necessary.
Training for intelligent exploration.    We believe training an agent not only to always exploit but to systematically gather information may be necessary to develop foundational exploration agents. In this regard, LLMs should be explored from a meta-reinforcement learning perspective. Nie et al. (2024) show first evidence that finetuning an agent on exploration trajectories can help in multi-armed bandits. Further, our results demonstrated that exploration capabilities may vary with prompt-engineering. Future agents should be less dependent on prompting and implicitly self-explore.
Memory architectures.    Many agent systems currently focus on developing memory architectures (Wang et al., 2024; Anokhin et al., 2024) that load the LLM context dynamically. Testing such architectures for systematic exploration, similarly to Lu et al. (2024), should be an important dimension in the evaluation. Further, building a meta memory that allows to reuse exploration-specific information across tasks and environments may be a fruitful path.
Small exploration models for real time applications.    We believe giving low parameter models the ability to explore is a necessary steps toward intelligent embedded and real time agents. As larger LLMs have shown to be more proficient at exploration, distillation methods could possibly be developed to teach exploration to smaller models, running on more devices.
Improving sequence understanding. During our experiments, we found that agents increasingly produce invalid actions and misunderstand longer sequences of past actions. Developing agents specifically made for long-sequence understanding is crucial for future LLM-based agents. It would be interesting to see if LLMs finetuned on sequential data (Kim et al., 2024; Zeng et al., 2023; Song et al., 2024) also exhibit better exploration capabilities.
6 Conclusion
In summary, with this work, we proposed a decomposition based on the optimal exploitation return to evaluate LLM-based agents for intelligent state-space exploration. Our evaluation includes an adaption of a popular environment from the reinforcement learning domain that goes beyond stateless multi-armed bandits and is less domain biased than other popular benchmarks. We believe developing systems that can do both, intelligent exploitation and exploration, are a crucial step towards truly generalist agents. Future work could apply our framework to evaluate agents in continuous and more complex environments. For example, we plan to include more randomization, stochasticity, and task-specific exploration. A limitation of our domain is that rewards are randomly spread and high return, therefore, correlates with high state-space coverage. By developing environments that combine complex exploration with underlying rules, domain knowledge and different tasks, we could use our framework to test more intelligent exploration of agents. We believe, as exploitation and exploration are both difficult problems, it is essential to evaluate both in isolation.
Appendix A Environment
In this section, we provide further information about the environments used in in the main paper.
A.1 Layouts
Figure 4: 4x4 grid variations.
Figure 5: 5x5 grid variations.
Figure 6: 7x7 grid variations.
A.2 Prompt Template
Large Language Model Agent Your past episodes: <HISTORY> You are controlling an agent in an unknown world. Over a total of 20 episodes, you can interact with objects in the environment. <INSTRUCTION> You have <MAX STEPS> door interactions per episode but can pick up three balls, keys, or boxes. Current episode: <EPISODIC HISTORY> Which object do you want to interact with next? Reply with one object enclosed with <and >, e.g. <door>. What is your next action?
Figure 7: Prompt Template
A.3 Abstraction Wrapper
Our room worlds are built upon the minigrid framework. To address the limitations of the low-level action and observation spaces, we use an abstraction wrapper function. This function transforms raw observations and actions into higher-level abstractions, allowing agents to reason more effectively about interactions with objects and doors within a room. By abstracting low-level details, the evaluation focuses on the agent’s exploration strategy rather than basic navigation capabilities. The abstraction wrapper defines the state as a discrete set of available objects and doors within the current room. Each room is uniquely configured to ensure that the agent can differentiate between them. Agent actions are simplified to interactions with specific objects or doors (e.g., choosing to open a door or collect an object). Once the agent selects an action at this abstract level, the wrapper translates it into a sequence of low-level actions using a predefined low-level policy to execute the selected behavior. Our wrapper can be seen as bridging the minigrid framework to a representation similar to text-environments like TextWorld.
A.4 Optimal Exploitation
We formulate the computation of the optimal exploitation return $R^{opt}$ as an instance of the Orienteering Problem. We represent the explored balls in the state-space of history $h_{t}$ as a graph $G=(V,E)$, where $V$ denotes the set of balls and $E$ the paths between them. Each ball $i\in V$ is assigned a reward $p_{i}$, and each edge $(i,j)\in E$ has an associated travel cost $c_{ij}$. Starting at a designated initial state $s\in V$, the goal is to determine a path $T$ through a subset of nodes $S\subseteq V$ that maximizes the cumulative rewards $\sum_{i\in S}p_{i}$, while ensuring that the total travel cost $\sum_{(i,j)\in T}c_{ij}$ does not exceed a predefined budget $C_{\text{max}}$. Formally, this optimization problem is expressed as:
$\max_{x_{ij},y_{i}}$ $\sum_{i\in V}p_{i}y_{i}$ (5a)
subject to $\sum_{(i,j)\in E}c_{ij}x_{ij}\leq C_{\text{max}},$ (5b)
$\sum_{j\in V}x_{ij}-\sum_{j\in V}x_{ji}=\begin{cases}1&\text{if }i=s,\\ 0&\text{otherwise},\end{cases}\quad\forall i\in V,$ (5c)
$x_{ij}\in\{0,1\},\quad\forall(i,j)\in E,$ (5d)
$y_{i}\in\{0,1\},\quad\forall i\in V.$ (5e)
where $x_{ij}$ and $y_{i}$ are binary decision variables indicating whether edge $(i,j)$ and ball $i$ is visited, respectively. Exact solvers, such as integer linear programming methods, or heuristic algorithms can be used to solve the optimization problem. We use a brute force approach. Note that our framework is not bound to any specific exploitation method.
A.5 Exemplary Environment Interaction
Targeted-Exploration Prompting Your past episodes: Episode 1: You see: dodger_blue door, tangerine door >dodger_blue door Reward: 0 You see: apricot door, dodger_blue door, cerulean door, honeydew door >cerulean door Reward: 0 You see: rosewood ball, teal door, cerulean door >rosewood ball Reward: 3 You see: teal door, cerulean door >teal door Reward: 0 You see: turquoise ball, magenta door, teal door, khaki ball >turquoise ball Reward: 2 You see: magenta door, teal door, khaki ball >khaki ball Reward: 3 You see: magenta door, teal door You are controlling an agent in an unknown world. Over a total of 20 episodes, you can interact with objects in the environment. You have 8 door interactions per episode but can pick up three balls, keys, or boxes. Based the current and past trials, explore the environment to collect information that may help to become better at maximizing the reward. Current episode: You see: dodger_blue door, tangerine door >tangerine door Reward: 0 You see: midnight_blue ball, tangerine door >midnight_blue ball Reward: 5 You see: tangerine door >tangerine door Reward: 0 You see: tangerine door, dodger_blue door Which object do you want to interact with next? Reply with one object enclosed with <and >,e.g. <door>. What is your next action?
Figure 8: Prompt Template
Appendix B Additional Results
B.1 Measuring Exploration with Large Language Model Exploitation
In the following, we build an agent, in which Llama 3.3 conducts the exploitation of the current history. We tested if LLMs, similarly to optimal returns, can find reliable exploitation steps from a history of trajectories. We found in figure 9 that LLMs struggle to achieve reliable and differentiable exploitation returns. These additional findings show that there is also a large research gap when it comes to understanding multiple sequential trajectories and finding the optimal sequence of actions with LLMs. We exemplary illustrate and hypothesize that, when exploration becomes more demanding, then LLM exploitation returns are similarly not reliable. For a proper assessment of LLM exploration, it is necessary to rely on exact methods, as proposed in our paper.
(a) LLM Exploration Gap
(b) Optimal Exploration Gap (ours)
Figure 9: We find that LLM exploitation is similarly hard to exploration and cannot reliably measure exploration progress. Instead, we propose to rely on exact methods to fairly evaluate exploration.
B.2 Learning Curves
(a) Agent Return
(b) Exploration Gap
(c) Agent Return
(d) Exploration Gap
Figure 10: Overall
(a) Agent Return
(b) Exploration Gap
(c) Agent Return
(d) Exploration Gap
Figure 11: 4x4 Grid
(a) Agent Return
(b) Exploration Gap
(c) Agent Return
(d) Exploration Gap
Figure 12: 5x5 Grid
(a) Agent Return
(b) Exploration Gap
(c) Agent Return
(d) Exploration Gap
Figure 13: 7x7 Grid
B.3 Total Gap Decomposition
Last Episode Mean
Total Gap Exploitation Gap Exploration Gap Total Gap Exploitation Gap Exploration Gap
<30B Mistral (7B) 0.80 0.25 (0.31) 0.55 (0.69) 0.78 0.20 (0.26) 0.57 (0.74)
Llama 3.1 (7B) 0.84 0.35 (0.41) 0.49 (0.59) 0.82 0.24 (0.28) 0.57 (0.72)
Gemma 2 (9B) 0.75 0.15 (0.17) 0.60 (0.83) 0.69 0.08 (0.10) 0.60 (0.90)
Mistral Small (22B) 0.68 0.28 (0.41) 0.40 (0.59) 0.69 0.20 (0.29) 0.49 (0.71)
Gemma 2 (27B) 0.57 0.18 (0.32) 0.38 (0.68) 0.52 0.12 (0.22) 0.40 (0.78)
>30B Llama 3.3 (70B) 0.50 0.23 (0.40) 0.27 (0.60) 0.50 0.18 (0.32) 0.32 (0.68)
Mistral Large (123B) 0.33 0.26 (0.75) 0.07 (0.25) 0.38 0.25 (0.62) 0.13 (0.38)
Closed GPT-4o 0.44 0.14 (0.31) 0.30 (0.69) 0.49 0.14 (0.27) 0.35 (0.73)
GPT-o1-mini 0.43 0.38 (0.79) 0.06 (0.21) 0.47 0.32 (0.66) 0.16 (0.34)
Gemini 1.5 Pro 0.52 0.30 (0.58) 0.22 (0.42) 0.53 0.23 (0.41) 0.30 (0.59)
Claude 3.5 Sonnet 0.38 0.09 (0.29) 0.29 (0.71) 0.41 0.09 (0.26) 0.32 (0.74)
Baseline Random Walk 0.85 0.53 (0.62) 0.32 (0.38) 0.86 0.37 (0.42) 0.49 (0.58)
Table 3: Total gap decomposition averaged over 4x4, 5x5, and 7x7 grids
Last Episode Mean
Total Gap Exploitation Gap Exploration Gap Total Gap Exploitation Gap Exploration Gap
<30B Mistral (7B) 0.84 0.50 (0.59) 0.34 (0.41) 0.80 0.41 (0.51) 0.39 (0.49)
Llama 3.1 (7B) 0.88 0.34 (0.36) 0.54 (0.64) 0.83 0.24 (0.26) 0.58 (0.74)
Gemma 2 (9B) 0.74 0.31 (0.35) 0.43 (0.65) 0.58 0.14 (0.20) 0.44 (0.80)
Mistral Small (22B) 0.64 0.30 (0.46) 0.33 (0.54) 0.64 0.24 (0.34) 0.40 (0.66)
Gemma 2 (27B) 0.57 0.24 (0.43) 0.33 (0.57) 0.52 0.16 (0.30) 0.36 (0.70)
>30B Llama 3.3 (70B) 0.53 0.25 (0.35) 0.28 (0.65) 0.53 0.18 (0.26) 0.34 (0.74)
Mistral Large (123B) 0.23 0.18 (0.79) 0.05 (0.21) 0.31 0.21 (0.68) 0.10 (0.32)
Closed GPT-4o 0.50 0.12 (0.20) 0.37 (0.80) 0.55 0.12 (0.20) 0.43 (0.80)
GPT-o1-mini 0.39 0.31 (0.77) 0.08 (0.23) 0.47 0.27 (0.59) 0.20 (0.41)
Gemini 1.5 Pro 0.40 0.30 (0.82) 0.10 (0.18) 0.51 0.27 (0.55) 0.23 (0.45)
Claude 3.5 Sonnet 0.35 0.10 (0.36) 0.25 (0.64) 0.37 0.09 (0.29) 0.28 (0.71)
Baseline Random Walk 0.83 0.55 (0.64) 0.28 (0.36) 0.86 0.39 (0.44) 0.47 (0.56)
Table 4: Total gap decomposition of the 4x4 grids
Last Episode Mean
Total Gap Exploitation Gap Exploration Gap Total Gap Exploitation Gap Exploration Gap
<30B Mistral (7B) 0.75 0.16 (0.23) 0.60 (0.77) 0.73 0.12 (0.17) 0.62 (0.83)
Llama 3.1 (7B) 0.89 0.47 (0.53) 0.42 (0.47) 0.85 0.30 (0.35) 0.55 (0.65)
Gemma 2 (9B) 0.72 0.05 (0.05) 0.66 (0.95) 0.69 0.02 (0.02) 0.67 (0.98)
Mistral Small (22B) 0.72 0.36 (0.51) 0.36 (0.49) 0.70 0.23 (0.34) 0.47 (0.66)
Gemma 2 (27B) 0.45 0.07 (0.15) 0.38 (0.85) 0.46 0.07 (0.13) 0.40 (0.87)
>30B Llama 3.3 (70B) 0.48 0.21 (0.41) 0.27 (0.59) 0.47 0.16 (0.33) 0.31 (0.67)
Mistral Large (123B) 0.43 0.34 (0.79) 0.09 (0.21) 0.46 0.29 (0.61) 0.17 (0.39)
Closed GPT-4o 0.44 0.16 (0.33) 0.28 (0.67) 0.47 0.14 (0.26) 0.33 (0.74)
GPT-o1-mini 0.49 0.44 (0.77) 0.05 (0.23) 0.47 0.32 (0.65) 0.15 (0.35)
Gemini 1.5 Pro 0.60 0.30 (0.47) 0.30 (0.53) 0.57 0.20 (0.35) 0.37 (0.65)
Claude 3.5 Sonnet 0.45 0.08 (0.25) 0.36 (0.75) 0.47 0.09 (0.24) 0.38 (0.76)
Baseline Random Walk 0.82 0.50 (0.60) 0.32 (0.40) 0.85 0.36 (0.41) 0.50 (0.59)
Table 5: Total gap decomposition of the 5x5 grids
Last Episode Mean
Total Gap Exploitation Gap Exploration Gap Total Gap Exploitation Gap Exploration Gap
<30B Mistral (7B) 0.79 0.09 (0.10) 0.71 (0.90) 0.80 0.08 (0.10) 0.71 (0.90)
Llama 3.1 (7B) 0.75 0.25 (0.33) 0.50 (0.67) 0.78 0.19 (0.23) 0.59 (0.77)
Gemma 2 (9B) 0.79 0.10 (0.10) 0.70 (0.90) 0.79 0.09 (0.09) 0.70 (0.91)
Mistral Small (22B) 0.68 0.18 (0.26) 0.50 (0.74) 0.74 0.14 (0.18) 0.60 (0.82)
Gemma 2 (27B) 0.67 0.23 (0.36) 0.44 (0.64) 0.59 0.13 (0.24) 0.46 (0.76)
>30B Llama 3.3 (70B) 0.49 0.22 (0.45) 0.27 (0.55) 0.49 0.18 (0.36) 0.31 (0.64)
Mistral Large (123B) 0.34 0.27 (0.67) 0.07 (0.33) 0.37 0.25 (0.58) 0.12 (0.42)
Closed GPT-4o 0.38 0.12 (0.38) 0.26 (0.62) 0.47 0.15 (0.36) 0.32 (0.64)
GPT-o1-mini 0.43 0.38 (0.85) 0.04 (0.15) 0.48 0.35 (0.73) 0.13 (0.27)
Gemini 1.5 Pro 0.57 0.30 (0.45) 0.27 (0.55) 0.51 0.21 (0.33) 0.30 (0.67)
Claude 3.5 Sonnet 0.38 0.08 (0.20) 0.29 (0.80) 0.42 0.10 (0.21) 0.33 (0.79)
Baseline Random Walk 0.89 0.55 (0.61) 0.34 (0.39) 0.86 0.35 (0.40) 0.50 (0.60)
Table 6: Total return gap decomposition of the 7x7 grids
B.4 Other Statistics
Exploitation Return Agent Return Coverage Redundancy Sample Efficiency
<30B Mistral (7B) 9.47 4.13 38.88 0.92 13.82
Llama 3.1 (7B) 10.56 3.27 37.03 0.92 31.25
Gemma 2 (9B) 8.53 5.18 30.10 0.95 7.25
Mistral Small (22B) 12.54 6.67 46.19 0.91 35.44
Gemma 2 (27B) 12.82 8.99 49.99 0.92 14.13
>30B Llama 3.3 (70B) 14.94 10.32 56.93 0.91 19.31
Mistral Large (123B) 19.20 13.86 78.59 0.88 19.56
Closed GPT-4o 14.22 11.38 54.33 0.92 20.73
GPT-o1-mini 19.40 11.76 87.73 0.86 27.88
Gemini 1.5 Pro 16.09 9.91 67.02 0.89 23.25
Claude 3.5 Sonnet 15.03 13.08 58.00 0.91 12.08
Baseline Random Walk 14.18 3.13 69.11 0.89 51.25
Table 7: Other statistics averaged over 4x4, 5x5, and 7x7 grids
Exploitation Return Agent Return Coverage Redundancy Sample Efficiency
<30B Mistral (7B) 14.57 3.50 50.00 0.92 19.67
Llama 3.1 (7B) 10.30 2.77 44.39 0.91 21.50
Gemma 2 (9B) 12.83 5.70 45.45 0.94 9.33
Mistral Small (22B) 15.00 8.13 58.94 0.91 32.00
Gemma 2 (27B) 15.00 9.53 63.64 0.92 15.33
>30B Llama 3.3 (70B) 16.03 10.67 63.79 0.91 20.83
Mistral Large (123B) 21.33 17.27 83.48 0.89 18.83
Closed GPT-4o 14.07 11.27 60.61 0.92 19.00
GPT-o1-mini 20.67 13.87 92.73 0.88 33.33
Gemini 1.5 Pro 20.07 13.27 88.48 0.88 35.00
Claude 3.5 Sonnet 16.73 14.50 58.48 0.92 11.17
Baseline Random Walk 16.10 3.77 78.79 0.88 49.67
Table 8: Other statistics of the 4x4 grids
Exploitation Return Agent Return Coverage Redundancy Sample Efficiency
<30B Mistral (7B) 8.13 4.90 40.34 0.92 11.50
Llama 3.1 (7B) 11.53 2.20 41.84 0.93 39.67
Gemma 2 (9B) 6.83 5.83 22.99 0.96 7.33
Mistral Small (22B) 12.73 5.60 48.28 0.91 41.00
Gemma 2 (27B) 12.37 11.00 50.23 0.93 12.17
>30B Llama 3.3 (70B) 14.40 10.27 60.34 0.91 18.67
Mistral Large (123B) 17.93 11.30 86.90 0.87 26.67
Closed GPT-4o 14.13 11.03 58.62 0.91 18.83
GPT-o1-mini 18.73 10.20 94.94 0.86 31.00
Gemini 1.5 Pro 13.80 8.00 71.03 0.89 25.33
Claude 3.5 Sonnet 12.73 11.07 59.77 0.91 13.00
Baseline Random Walk 13.37 3.50 72.53 0.89 57.67
Table 9: Other statistics of the 5x5 grids
Exploitation Return Agent Return Coverage Redundancy Sample Efficiency
<30B Mistral (7B) 5.70 4.00 26.28 0.93 6.88
Llama 3.1 (7B) 9.83 4.83 24.87 0.93 23.83
Gemma 2 (9B) 5.93 4.00 21.86 0.95 6.67
Mistral Small (22B) 9.90 6.27 31.35 0.92 25.83
Gemma 2 (27B) 11.10 6.43 36.09 0.93 13.83
>30B Llama 3.3 (70B) 14.40 10.03 46.67 0.91 15.33
Mistral Large (123B) 18.33 13.00 65.38 0.87 16.50
Closed GPT-4o 14.53 12.20 39.49 0.92 21.33
GPT-o1-mini 18.80 11.20 75.51 0.85 25.33
Gemini 1.5 Pro 14.40 8.47 41.54 0.91 14.67
Claude 3.5 Sonnet 13.93 12.27 55.26 0.89 13.85
Baseline Random Walk 13.07 2.13 56.03 0.88 42.17
Table 10: Other statistics of the 7x7 grids
B.5 What Impact do Agent Instructions have on Exploration?
We demonstrate how our evaluation framework can be used to analyze the impact of prompt-engineering on exploration. All prompts used to evaluate the impact of prompt-engineering on exploration behaviors is illustrated in figure 14. To test the limits of LLM exploration, we design two prompts besides the task-oriented exploration (middle) from previous experiments. One of the prompts is written to analyze if agents can do full exploitation of their history (left), while the other should make agents explore aimlessly without mentioning a task.
Figure 14: Our prompts on a scale from optimal exploitation to exploration.
(a) Mistral (7B)
(b) Llama 3.1 (7B)
(c) Gemma 2 (9B)
(d) Mistral Small (22B)
(e) Gemma 2 (27B)
(f) Llama 3.3 (70B)
(g) Mistral Large (123B)
Figure 15: Impact of the agent instruction on open-source model behavior.
(a) GPT-4o
(b) o1-mini
(c) Gemini 1.5 Pro
(d) Claude 3.5 Sonnet
Figure 16: Impact of the prompt on closed model behavior
B.5.1 Soft Exploration Lower Bound
Last Episode Mean
Total Gap Exploitation Gap Exploration Gap Total Gap Exploitation Gap Exploration Gap
<30B Mistral (7B) 0.78 0.09 (0.10) 0.69 (0.90) 0.79 0.08 (0.09) 0.71 (0.91)
Llama 3.1 (7B) 0.88 0.25 (0.25) 0.63 (0.75) 0.81 0.13 (0.14) 0.68 (0.86)
Gemma 2 (9B) 0.66 0.00 (0.00) 0.66 (1.00) 0.72 0.06 (0.06) 0.66 (0.94)
Mistral Small (22B) 0.76 0.21 (0.26) 0.55 (0.74) 0.74 0.10 (0.13) 0.65 (0.87)
Gemma 2 (27B) 0.85 0.21 (0.22) 0.64 (0.78) 0.79 0.12 (0.14) 0.67 (0.86)
>30B Llama 3.3 (70B) 0.58 0.06 (0.08) 0.53 (0.92) 0.60 0.05 (0.07) 0.55 (0.93)
Mistral Large (123B) 0.40 0.06 (0.13) 0.34 (0.87) 0.44 0.04 (0.10) 0.40 (0.90)
Closed GPT-4o 0.65 0.04 (0.06) 0.61 (0.94) 0.66 0.03 (0.05) 0.63 (0.95)
o1-mini 0.51 0.43 (0.82) 0.08 (0.18) 0.50 0.28 (0.55) 0.21 (0.45)
Gemini 1.5 Pro 0.73 0.06 (0.07) 0.66 (0.93) 0.69 0.03 (0.03) 0.66 (0.97)
Claude 3.5 Sonnet 0.48 0.00 (0.00) 0.48 (1.00) 0.48 0.00 (0.00) 0.48 (1.00)
Baseline Random Walk 0.89 0.55 (0.61) 0.34 (0.39) 0.86 0.35 (0.40) 0.50 (0.60)
Table 11: Return gap decomposition of the 7x7 grids
Last Episode Mean
Total Gap Exploitation Gap Exploration Gap Total Gap Exploitation Gap Exploration Gap
<30B Mistral (7B) -2.10 1.23 (-0.85) -2.50 (0.09) -0.72 -1.31 (-5.21) -0.65 (0.56)
Llama 3.1 (7B) 16.93 -1.91 (-22.82) 26.43 (11.24) 4.71 -31.25 (-38.33) 16.40 (11.17)
Gemma 2 (9B) -16.80 -100.00 (-100.00) -5.26 (10.70) -8.35 -28.40 (-28.52) -5.82 (2.77)
Mistral Small (22B) 11.76 17.04 (-0.54) 9.83 (0.19) 0.78 -30.79 (-30.00) 8.03 (6.69)
Gemma 2 (27B) 25.70 -11.53 (-38.55) 45.72 (21.83) 34.50 -5.38 (-41.08) 45.76 (12.75)
>30B Llama 3.3 (70B) 19.50 -73.66 (-81.69) 96.81 (67.30) 22.00 -73.53 (-81.42) 78.20 (46.57)
Mistral Large (123B) 19.95 -76.79 (-80.06) 409.96 (163.94) 18.54 -82.24 (-82.47) 219.95 (115.75)
Closed GPT-4o 71.96 -66.32 (-84.95) 135.81 (51.97) 39.42 -77.77 (-86.81) 93.63 (48.09)
o1-mini 19.24 11.00 (-2.86) 92.31 (15.61) 2.88 -19.92 (-24.88) 65.47 (66.61)
Gemini 1.5 Pro 27.64 -79.59 (-84.88) 147.64 (69.16) 36.86 -86.76 (-91.17) 122.21 (44.49)
Claude 3.5 Sonnet 28.13 -100.00 (-100.00) 64.53 (24.60) 13.78 -100.00 (-100.00) 48.43 (27.18)
Table 12: Percentage change in the return gap decomposition of the 7x7 grids
Exploitation Return Agent Return Coverage Redundancy Sample Efficiency
<30B Mistral (7B) 6.07 4.33 21.86 0.95 10.43
Llama 3.1 (7B) 7.20 2.33 16.41 0.94 18.50
Gemma 2 (9B) 6.67 6.67 13.33 0.97 5.00
Mistral Small (22B) 8.90 4.70 26.47 0.94 27.67
Gemma 2 (27B) 7.13 3.00 20.51 0.95 14.33
>30B Llama 3.3 (70B) 9.30 8.13 26.22 0.95 12.17
Mistral Large (123B) 12.87 11.63 38.08 0.92 21.60
Closed GPT-4o 7.73 6.93 25.64 0.95 12.67
GPT-o1-mini 18.00 9.60 62.56 0.88 36.33
Gemini 1.5 Pro 6.60 5.40 14.10 0.96 5.00
Claude 3.5 Sonnet 10.20 10.20 23.59 0.96 5.00
Baseline Random Walk 13.07 2.13 56.03 0.88 42.17
Table 13: Other statistics of the 7x7 grids
Exploitation Return Agent Return Coverage Redundancy Sample Efficiency
<30B Mistral (7B) 6.43 8.33 -16.83 2.86 51.78
Llama 3.1 (7B) -26.78 -51.72 -34.02 0.61 -22.38
Gemma 2 (9B) 12.36 66.67 -39.00 1.95 -25.00
Mistral Small (22B) -10.10 -25.00 -15.54 1.81 7.10
Gemma 2 (27B) -35.74 -53.37 -43.16 2.47 3.61
>30B Llama 3.3 (70B) -35.42 -18.94 -43.82 4.01 -20.65
Mistral Large (123B) -29.82 -10.51 -41.76 5.94 30.91
Closed GPT-4o -46.79 -43.17 -35.06 3.15 -40.63
o1-mini -4.26 -14.29 -17.15 3.11 43.42
Gemini 1.5 Pro -54.17 -36.22 -66.05 5.78 -65.91
Claude 3.5 Sonnet -26.79 -16.85 -57.31 7.00 -63.89
Table 14: Percentage change in the other statistics of the 7x7 grids
B.5.2 Soft Exploration Upper Bound
Last Episode Mean
Total Gap Exploitation Gap Exploration Gap Total Gap Exploitation Gap Exploration Gap
<30B Mistral (7B) 0.79 0.21 (0.24) 0.57 (0.76) 0.78 0.18 (0.20) 0.60 (0.80)
Llama 3.1 (7B) 0.78 0.45 (0.55) 0.33 (0.45) 0.75 0.35 (0.44) 0.40 (0.56)
Gemma 2 (9B) 0.64 0.38 (0.58) 0.26 (0.42) 0.61 0.27 (0.43) 0.34 (0.57)
Mistral Small (22B) 0.68 0.57 (0.85) 0.11 (0.15) 0.63 0.45 (0.70) 0.18 (0.30)
Gemma 2 (27B) 0.48 0.45 (0.91) 0.03 (0.09) 0.52 0.41 (0.74) 0.12 (0.26)
>30B Llama 3.3 (70B) 0.72 0.63 (0.88) 0.09 (0.12) 0.71 0.50 (0.71) 0.21 (0.29)
Mistral Large (123B) 0.82 0.81 (0.99) 0.01 (0.01) 0.81 0.69 (0.83) 0.12 (0.17)
Closed GPT-4o 0.92 0.86 (0.93) 0.06 (0.07) 0.80 0.56 (0.69) 0.24 (0.31)
o1-mini 0.65 0.59 (0.84) 0.06 (0.16) 0.70 0.48 (0.69) 0.22 (0.31)
Gemini 1.5 Pro 0.96 0.91 (0.94) 0.05 (0.06) 0.85 0.75 (0.84) 0.11 (0.16)
Claude 3.5 Sonnet 0.70 0.70 (1.00) 0.00 (0.00) 0.76 0.71 (0.91) 0.05 (0.09)
Baseline Random Walk 0.89 0.55 (0.61) 0.34 (0.39) 0.86 0.35 (0.40) 0.50 (0.60)
Table 15: Return gap decomposition of the 7x7 grids
Last Episode Mean
Total Gap Exploitation Gap Exploration Gap Total Gap Exploitation Gap Exploration Gap
<30B Mistral (7B) -0.87 148.21 (139.65) -18.89 (-15.52) -2.38 107.55 (104.17) -15.41 (-11.24)
Llama 3.1 (7B) 3.40 77.68 (65.65) -34.03 (-32.32) -3.47 83.17 (93.99) -31.63 (-27.38)
Gemma 2 (9B) -19.31 295.83 (499.07) -63.01 (-53.41) -22.40 207.05 (386.99) -51.30 (-37.56)
Mistral Small (22B) -0.37 212.75 (226.00) -78.14 (-79.81) -14.77 226.15 (281.42) -70.10 (-62.79)
Gemma 2 (27B) -28.18 91.27 (151.68) -92.38 (-85.89) -10.56 215.73 (212.76) -74.49 (-66.01)
>30B Llama 3.3 (70B) 46.52 182.46 (95.38) -66.28 (-78.57) 45.61 177.53 (95.43) -32.02 (-54.58)
Mistral Large (123B) 144.41 201.33 (46.88) -85.06 (-96.00) 117.56 177.52 (41.58) -2.29 (-58.36)
Closed GPT-4o 143.82 623.63 (144.12) -77.76 (-88.18) 68.45 272.84 (94.56) -26.10 (-52.38)
o1-mini 51.48 53.81 (-0.33) 30.77 (1.79) 46.32 37.07 (-5.22) 71.72 (13.97)
Gemini 1.5 Pro 69.58 204.44 (109.78) -81.35 (-89.45) 69.06 262.88 (154.71) -64.75 (-75.49)
Claude 3.5 Sonnet 85.84 740.00 (406.57) -100.00 (-100.00) 78.58 614.58 (325.05) -84.62 (-88.36)
Table 16: Percentage change in the return gap decomposition of the 7x7 grids
Exploitation Return Agent Return Coverage Redundancy Sample Efficiency
<30B Mistral (7B) 8.36 4.14 35.16 0.92 15.95
Llama 3.1 (7B) 13.17 4.30 34.49 0.86 22.33
Gemma 2 (9B) 14.70 7.07 52.76 0.89 26.83
Mistral Small (22B) 17.50 6.30 57.69 0.87 25.50
Gemma 2 (27B) 19.00 10.13 71.15 0.85 25.67
>30B Llama 3.3 (70B) 17.87 5.53 81.60 0.83 30.33
Mistral Large (123B) 19.47 3.50 90.83 0.81 32.83
Closed GPT-4o 18.53 1.60 81.15 0.83 38.33
o1-mini 18.53 6.87 82.05 0.83 40.00
Gemini 1.5 Pro 18.67 0.73 78.46 0.73 19.67
Claude 3.5 Sonnet 20.00 6.00 86.54 0.83 15.00
Baseline Random Walk 13.07 2.13 56.03 0.88 42.17
Table 17: Other statistics of the 7x7 grids
Exploitation Return Agent Return Coverage Redundancy Sample Efficiency
<30B Mistral (7B) 46.62 3.57 33.80 -0.93 132.03
Llama 3.1 (7B) 33.90 -11.03 38.66 -7.99 -6.29
Gemma 2 (9B) 147.75 76.67 141.35 -6.31 302.50
Mistral Small (22B) 76.77 0.53 84.05 -4.83 -1.29
Gemma 2 (27B) 71.17 57.51 97.16 -8.19 85.54
>30B Llama 3.3 (70B) 24.07 -44.85 74.86 -8.28 97.83
Mistral Large (123B) 6.18 -73.08 38.92 -6.69 98.99
Closed GPT-4o 27.52 -86.89 105.52 -9.42 79.69
o1-mini -1.42 -38.69 8.66 -1.91 57.89
Gemini 1.5 Pro 29.63 -91.34 88.89 -19.41 34.09
Claude 3.5 Sonnet 43.54 -51.09 56.61 -7.52 8.33
Table 18: Percentage change in the other statistics of the 7x7 grids